OpenStudy (anonymous):

Graph the ellipse and locate the foci x^2/36+y^2/11=1. I got tthe following: 1/36 x^2 + (0)^2/11=1 1/35x^2=1 x^2=1x36 x=+- squar root of 36 x= +6, -6 ------------------- 0/36+1/11y^2=1 y^2=11 y= +square root of 11, - square root 11 --------------------------- answer is foci at (square root 11, 0) and (negatie square root 11,0)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay, what went wrong?

OpenStudy (anonymous):

your foci are wrong

OpenStudy (anonymous):

do you know how this is oriented?

OpenStudy (anonymous):



OpenStudy (anonymous):

Well, I know because the denominator of x^2 is grater than y^2, the major axis is horizontal. The other possible horizontal answer would be focie at (5,0) and (-5,0).

OpenStudy (anonymous):

But where did I go wrong when I got the other answer?

OpenStudy (anonymous):

let me ask again you need to know which way the ellipse is oriented to get the right answer

OpenStudy (anonymous):

The ellipse would be horizontal. .

OpenStudy (anonymous):

oh i see you answered, the major axis is horizontal

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

and the center is \((0,0)\) so the foci are to the left and right of the center

OpenStudy (anonymous):

That's is correct.

OpenStudy (anonymous):

you have \(a^2=36,b^2=11\) you need \(a^2-b^2\)

OpenStudy (anonymous):

Hold on. I am doing this.

OpenStudy (anonymous):

1296-121=1175

OpenStudy (anonymous):

ooh no don't square them ! they are already squared

OpenStudy (anonymous):

\[a^2=36,b^2=11,a^2-b^2=36-11=25\]

OpenStudy (anonymous):

Sorry, my bad.. I see a square and I square it.

OpenStudy (anonymous):

right, don't do tha? that makes \(c^2=25\) and so \(c=5\) making your foci 5 units to the left and right of the center

OpenStudy (anonymous):

So then that would make a 5 and a -5.

OpenStudy (anonymous):

Answer would be foci at (5,0) and (-5,0).

OpenStudy (anonymous):

I have been at this most of the day and have accomplished 8 questions. The book is not really explicit in explanations, so am trying my best. I am only on to see if I got it right or where I went wrong. Thank you for explanations. That makes everything much easier.

OpenStudy (anonymous):

Thanks everyone for your help tonight. Two questions out of eight were uncertain and they are now answered. Thank you so much for everyone's help. Will be back for sure.

OpenStudy (anonymous):

oh sorry i stepped away yes \((-5,0)\) and \((5,0)\)

OpenStudy (anonymous):

if you want to see nice easy worked out examples, google "ellipse purplemath" and they have many easy to follow